OpenStudy (anonymous):

Solve, using the substitution method. y = 2x + 6 2x - y = 6 (Points : 1) (14, 12) (12, 30) There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

There are infinite number of solutions

OpenStudy (lilg132):

2x - 2x + 6 = 6 2x - 2x = 6 - 6 0 = 0 infinite number of soloutions